Problems Treasury Bonds ASKED YLD TO MATURITY MATURITY COUPON BID ASKED CHG. 2019 SEP 15 4.875 ??? 105:4685 0.0938

1.) Assume that this is the Wall Street Journal listing for the trading day June 15, 2005. For the above T-Bond, calculate the ask yield.

2.) If the tax rate were 40%, what would be the coupon rate for an equivalent corporate bond?

3.) If the difference between what broker receives when she sells the bond and when she buys the bond is $3.13, what is the bid quote?

(I'm not understanding on how to calculate these problems. There were two formulas given, but I'm not sure which questions I'm suppose to calculate them with.)

rm = r (1 - t)

r = rm ÷ (1 - t)
Answer #1
1..Assuming the bond face value of $ 100 &
Annual coupons,
PV=Asked price=105.4685
$ Coupon amt.= 4.875
n=No.of periods to maturity June15,2005-Sep 15 ,2019= 14.5 annual periods
Using the formula to find PV of abond,
PV=($ coupon amt.*(1-(1+Yield)^-n)/Yield)*(Face Value/(1+Yield)^n)
105.4685=(4.875*(1-(1+r)^-14.5)/r)+(100/(1+r)^14.5
Solving for r, we get the ask-yield as
4.358
ie. 4.358%
2. If the tax rate were 40%
Coupon rate for an equivalent corporate bond=
4.875/(1-40%)*1=
8.125
ie. 8.125%
3..Bid Quote=105.4685-3.13=
102.3385